1. What gauge pressure in Pa must a pump produce in order to pump water from the bottom of Grand Canyon (730 m elevation) to the Indian gardens (1370 m)?

2. Calculate the pressure at the bottom of the container with 1.2 m high layer of oil (SG=0.82), 0.5 m layer of water and 0.3 m layer of mercury (SG=13.6)

3. A cube having a side 8 cm in length has a density of 900 kg/m³. What is the mass and unit weight of the cube?

4. Georgie is trying to float his toy boat on water during one rainy day. If it weighs 130 N and has a density of 9 g/cm³, what is the total volume of the toy and the buoyant force exerted on it?
1
Expert's answer
2020-05-05T03:47:08-0400

Solution.

1."h_1=730m;"

"h_2=1370m;"

"p=\rho g\Delta h; \Delta h=h_2-h_1;"

"\Delta h=1370m-730m=640m;"

"p=1000kg/m^3\sdot9.8N/kg\sdot640m=6.272\sdot10^6Pa;"

2."h_o=1.2m; \rho_o=0.82g/cm^3=820kg/m^3;"

"h_w=0.5m; \rho_w=1000kg/m^3;"

"h_m=0.3m; \rho_w=13.6g/cm^3=13600kg/m^3;"

"p=\rho gh;"

"p_o=820kg/m^3\sdot 9.8N/kg\sdot1.2m=9643.2Pa=9.643kPa;"

"p_w=1000kg/m^3\sdot9.8N/kg\sdot0.5m=4900Pa=4.9kPa;"

"p_m=13600kg/m^3\sdot9.8N/kg\sdot0.3m=39984Pa=39.984kPa;"

"p=p_o+p_w+p_m;"

"p=9.643kPa+4.9kPa+39.984kPa=54.527kPa;"

3."a=8cm=0.08m;"

"\rho=900kg/m^3;"

"m=\rho\sdot V;"

"V=a^3;" "V=(0.08m)^3=0.000512m^3;"

"m=900kg/m^3\sdot0.000512m^3=0.4608kg;"

"P=mg; P=0.4608kg\sdot9.8N/kg=4.52N;"

4."P=130N;"

"\rho=9g/cm^3=9000kg/m^3;"

"m=\rho V \implies V=\dfrac{m}{\rho};"

"P=mg\implies m=\dfrac{P}{g};"

"m=\dfrac{130N}{9.8N/kg}=13.265kg;"

"V=\dfrac{13.265kg}{9000kg/m^3}=0.00147m^3;"

"F_A=\rho_w gV;"

"F_A=1000kg/m^3\sdot9.8N/kg\sdot0.00147m^3=14.4N;"

Answer: 1. "p=6.272\sdot10^6Pa;"

2."p=54.527kPa;"

3."m=0.4608kg;" "P=4.52N;"

4."V=0.00147m^3;" "F_A=14.4N;"